# HiCMAE: Hierarchical Contrastive Masked Autoencoder for Self-Supervised Audio-Visual Emotion Recognition

### v1

@@ 3.1.2 Encoder
Modality-Specific Encoder. As a general self-supervised pre-training framework, there can be many choices for the architecture of modality-specific encoders. For simplicity, we follow MAE he2022masked and VideoMAE tong2022videomae to employ standard Transformer vaswani2017attention as audio and video encoders. Other efficient architectures such as LGI-Former sun2023mae can be explored in future work. The audio and video encoders consist of $N_{s}$ Transformer layers. Each Transformer layer is mainly composed of Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN):
@@
where $\mathbf{E}^{0}_{m}=\mathbf{X}_{m}^{{}^{\prime\prime}}$, $m\in\{a,v\}$ denotes the audio or video modality, $j\in\{1,...,N_{s}\}$ is the layer index, and LN stands for layer normalization vaswani2017attention . For MHSA in Eq. (1), it calculates dot-product attention to learn intrinsic dependency relationships in input audio/video tokens:
@@
where $\mathbf{W}^{Q}_{h}\in\mathbb{R}^{C\times d_{h}}$, $\mathbf{W}^{K}_{h}\in\mathbb{R}^{C\times d_{h}}$, $\mathbf{W}^{V}_{h}\in\mathbb{R}^{C\times d_{h}}$, $\mathbf{W}^{O}\in\mathbb{R}^{C\times C}$, $H$ is the number of attention heads, and $d_{h}=C/H$ is the feature dimension in each attention head. For FFN in Eq. (1), it comprises two linear projection layers with a GELU hendrycks2016gaussian activation function in between:
@@
Cross-Modal Fusion Encoder. After unimodal feature encoding, HiCMAE employs a cross-modal fusion encoder to capture meaningful interactions in audio-visual modalities and enable cross-modal reinforcement tsai2019multimodal ; sun2023efficient . We mainly utilize multi-head cross-attention (MHCA) to implement this module. MHCA shares the same spirit with MHSA. The main difference is that MHCA accepts two modalities as input, with one as the target modality and the other as the source modality. The formulation of MHCA is given as follows:
@@ 3.1.3 Decoder
After feature encoding, HiCMAE utilizes two modality-specific decoders for final masked audio-visual reconstruction. Previous studies have demonstrated that a small-capacity model is enough to reconstruct masked information he2022masked ; tong2022videomae ; huang2022masked . Therefore, we follow them to employ a lightweight Transformer-based model as the decoder, whose number of layers is much less than that of the encoder, to largely reduce the computation burden during self-supervised pre-training.
@@
Hierarchical Skip Connections. As shown in Fig. 3, unlike previous audio-visual masked autoencoders which solely operate on encoder representation from the last layer and neglect explicit guidance on other layers gong2023contrastive ; georgescu2023audiovisual ; huang2023mavil ; sadok2023vector , HiCMAE incorporates hierarchical skip connections between intermediate encoder and decoder layers to explicitly steer encoder feature learning of different levels and assist the decoder to complete the task of masked audio-visual reconstruction. Specifically, HiCMAE adds an MHCA layer before each (except for the first) Transformer layer in the decoder. Given the tokens $\mathbf{D}^{k}_{m}$ ($m\in\{a,v\}$) output by the $k$th Transformer layer, the skip connection constructed by MHCA enables it to directly attend to the intermediate encoder representations $\mathbf{E}^{j}_{m}$ with different levels of information (from high-level coarse-grained to low-level fine-grained) and extract useful information to reinforce itself. Then, as usual, a standard Transformer layer is followed to exploit global self-attention to infer masked information. The whole process can be formulated as follows:
@@
where $\mathbf{X}_{m}$ ($m\in\{a,v\}$) denotes the original input, $\hat{\mathbf{X}}_{m}$ is the reconstructed data, $\mathcal{M}_{m}$ denotes the masked positions. Note that, for the visual modality, we follow sun2023mae to reconstruct both spatial appearance and temporal motion information (i.e., frame difference signals).
@@ 3.2 Hierarchical Cross-Modal Contrastive Learning
To align with hierarchical skip connections in Section 3.1.3, we use the same $N_{c}$ selected encoder layers ($i.e.,\{q_{1},...,q_{N_{c}}\}$, $q_{i}\in\{1,...,N_{s}\}$) for HCMCL. Since HCMCL is conducted within a batch of samples, for convenience, we add an additional subscript $i$ to the notation of encoder tokens (i.e., $\mathbf{E}^{j}_{i,m}$) to indicate the sample index in this batch. Given a batch of $N$ audio tokens $\{\mathbf{E}^{j}_{1,a},...,\mathbf{E}^{j}_{N,a}\}$ and $N$ video tokens $\{\mathbf{E}^{j}_{1,v},...,\mathbf{E}^{j}_{N,v}\}$ from $j$th ($j\in\{q_{1},...,q_{N_{c}}\}$) encoder layer, we first perform global average pooling to obtain the sample-level features, i.e., $\mathbf{e}^{j}_{i,a}=\textrm{AvgPool}(\mathbf{E}^{j}_{i,a})$ and $\mathbf{e}^{j}_{i,v}=\textrm{AvgPool}(\mathbf{E}^{j}_{i,v})$. After that, we utilize symmetric InfoNCE loss oord2018representation for HCMCL. For a batch of audio features $\mathbf{e}^{j}_{a}=\{\mathbf{e}^{j}_{1,a},...,\mathbf{e}^{j}_{N,a}\}$ and video features $\mathbf{e}^{j}_{v}=\{\mathbf{e}^{j}_{1,v},...,\mathbf{e}^{j}_{N,v}\}$, the symmetric InfoNCE loss maximizes the cosine similarity of $N$ paired audio-visual features (i.e., from the same sample) in the batch while minimizing the cosine similarity of features of $N(N-1)$ incorrect pairings (i.e., from different samples):
@@ 3.4 Hierarchical Feature Fusion for Downstream Fine-tuning
After self-supervised pre-training, we discard the decoder and only use the encoder in downstream fine-tuning. To benefit downstream tasks, we utilize both cross-modal and unimodal features from the encoder. The cross-modal features come from the cross-modal fusion encoder. We simply perform global average pooling to output tokens from the last fusion layer. The unimodal features come from modality-specific encoders. To fully exploit features of different levels, we use learnable weights to combine features from different audio/video encoder layers. Thus, the overall feature can be obtained as follows:
@@
After obtaining the overall feature $\mathbf{e}$, we simply use a linear layer to project it to get the final prediction. For the classification task, we utilize the classic cross-entropy loss for model fine-tuning:
$\mathcal{L}_{\textrm{CLS}}=-\sum_{c=1}^{C}{y_{c}\log{\hat{y}_{c}}},$ (13)
where $\hat{\mathbf{y}}=[\hat{y}_{1},...,\hat{y}_{C}]$ denotes the prediction, $\mathbf{y}=[y_{1},...,y_{C}]$ is the target, and $C$ is the number of classes. For the regression task, we compute the mean square error between the target and the prediction:
@@ 4.1 Implementation Details
We pre-train HiCMAE on a very large audio-visual dataset VoxCeleb2 chung2018voxceleb2 . It contains more than one million video clips from over six thousand celebrities. VoxCeleb2 is split into a development set and a test set. In this paper, we only use its development set which has 1,092,009 video clips for self-supervised pre-training. For each video clip, we sample 16 consecutive frames with a temporal stride of 4 frames and follow sun2023mae to extract $160\times 160$ patch in each frame to obtain the video input with a size of $16\times 160\times 160\times 3$. The corresponding audio waveform (2.56s) is also extracted and converted into a 128-dimensional log Mel filterbank feature sequence using a 25ms Hanning window with a hop length of 10ms gong2023contrastive , resulting in the audio spectrogram with a size of $256\times 128$. The loss weight in Eq. (11) is set to $\lambda=0.0025$. The temperature factor for contrastive learning is fixed to $\tau=0.07$. We pre-train HiCMAE for 100 epochs using four Nvidia Tesla V100 GPUs with a total batch size of 160 and a base learning rate of $43e-$ 4. It takes about five days to complete the pre-training. For downstream tasks, we fine-tune the pre-trained model for 50 or 100 epochs with a total batch size of 56 and a base learning rate of $31e-$ 3. During inference, we follow sun2023mae to uniformly sample two clips from each video and calculate their mean score as the final prediction. Other hyper-parameters for pre-training and fine-tuning can refer to tong2022videomae ; huang2022masked ; gong2023contrastive for details.
@@ 4.2.1 Datasets
MAFW liu2022mafw is a large-scale multimodal compound in-the-wild affective dataset. It consists of 10,045 video clips annotated with 11 common emotions (including seven basic emotions, contempt, anxiety, helplessness, and disappointment). Each video clip is also accompanied by several textual sentences to describe the subject’s affective behaviors. The dataset provides an 11-class single-labeled set (9,172 video clips) and a 43-class compound set (8,996 video clips). For model evaluation, we follow the original paper to adopt a 5-fold cross-validation protocol.
DFEW jiang2020dfew comprises 16,372 video clips which are extracted from over 1,500 high-definition movies. This dataset presents several challenging characteristics, such as extreme illumination and occlusion. The video clips are annotated with seven basic emotions (i.e., happy, sad, neutral, anger, surprise, disgust, and fear). To align with previous work jiang2020dfew ; sun2023mae ; zhang2023transformer , we perform 5-fold cross-validation on 11,697 single-labeled clips for evaluation.
MER-MULTI lian2023mer provides 3,373 training video clips originating from Chinese TV series and movies. The dataset is annotated with six emotions, including neutral, anger, happiness, sadness, worry, and surprise. We follow the original paper to conduct 5-fold cross-validation on 3,373 video clips for hyperparameter tuning and evaluate the model on a held-out test set with 411 video clips.
@@ 4.2.2 Results on MAFW
Wav2Vec2.0 baevski2020wav2vec A 95 18 59.01 9.39 26.08 31.47 32.04 46.52 9.91 1.69 12.23 3.05 6.04 21.59 29.69
HuBERT hsu2021hubert A 95 18 54.97 15.49 31.20 28.64 36.88 58.39 12.52 2.54 12.55 5.34 16.48 25.00 32.60
WavLM-Plus chen2022wavlm A 95 18 55.62 17.21 40.48 36.65 36.53 57.44 11.12 2.12 11.35 9.54 11.54 26.33 34.07
@@
ResNet-18 he2016deep $\times$ V 11 - 45.02 9.25 22.51 70.69 35.94 52.25 39.04 0.00 6.67 0.00 0.00 25.58 36.65
ViT dosovitskiy2020image $\times$ V - - 46.03 18.18 27.49 76.89 50.70 68.19 45.13 1.27 18.93 1.53 1.65 32.36 45.04
C3D tran2015learning $\times$ V 78 39 51.47 10.66 24.66 70.64 43.81 55.04 46.61 1.68 24.34 5.73 4.93 31.17 42.25
ResNet-18+LSTM liu2022mafw $\times$ V - - 46.25 4.70 25.56 68.92 44.99 51.91 45.88 1.69 15.75 1.53 1.65 28.08 39.38
ViT+LSTM liu2022mafw $\times$ V - - 42.42 14.58 35.69 76.25 54.48 68.87 41.01 0.00 24.40 0.00 1.65 32.67 45.56
C3D+LSTM liu2022mafw $\times$ V - - 54.91 0.47 9.00 73.43 41.39 64.92 58.43 0.00 24.62 0.00 0.00 29.75 43.76
Former-DFER zhao2021former $\times$ V 18 9 58.23 11.45 31.29 75.06 43.07 63.81 46.02 0.42 26.22 2.88 2.25 32.79 45.31
T-ESFL liu2022mafw $\times$ V - - 62.70 2.51 29.90 83.82 61.16 67.98 48.50 0.00 9.52 0.00 0.00 33.28 48.18
T-MEP zhang2023transformer $\times$ V 5 6 52.91 17.41 28.01 80.79 49.42 58.73 49.54 0.00 26.18 2.25 3.56 33.53 47.53
DFER-CLIP zhao2023dferclip V 153 92 - - - - - - - - - - - 39.89 52.55
SVFAP sun2023svfap V 78 44 64.60 25.20 35.68 82.77 57.12 70.41 58.58 8.05 32.42 8.40 9.89 41.19 54.28
MAE-DFER sun2023mae V 85 50 67.77 25.35 34.88 77.13 58.26 71.09 57.46 8.90 33.08 11.83 12.09 41.62 54.31
@@
ResNet-18+LSTM liu2022mafw $\times$ A+V - - 54.47 11.89 7.07 82.73 54.85 55.06 39.35 0.00 15.99 0.39 0.00 29.26 42.69
C3D+LSTM liu2022mafw $\times$ A+V - - 62.47 3.17 15.74 77.30 42.20 65.30 42.67 0.00 19.14 0.00 0.00 30.47 44.15
AMH yoon2020attentive $\times$ A+V - - 51.73 18.68 28.13 79.14 52.55 52.26 46.29 0.26 29.62 1.74 2.39 32.98 48.83
T-ESFL liu2022mafw $\times$ A+V - - 60.73 1.26 21.40 80.31 58.24 75.31 53.23 0.00 14.93 0.00 0.00 33.35 48.70
T-MEP* zhang2023transformer $\times$ A+V 61 22 54.98 22.11 32.23 82.79 50.90 62.50 49.93 0.87 29.27 8.09 6.70 36.40 48.17
T-MEP zhang2023transformer $\times$ A+V 61 22 57.04 24.85 36.09 78.96 50.83 61.85 51.28 1.29 38.47 6.46 1.70 37.17 51.15
@@
ResNet18+MDRE yoon2018multimodal $\times$ A+V+T - - 45.59 9.35 24.30 76.31 51.10 74.87 28.82 2.08 30.99 0.00 0.00 31.22 48.33
AMH yoon2020attentive $\times$ A+V+T - - 54.91 19.41 30.01 82.79 51.42 60.73 51.54 0.00 28.18 0.00 0.00 34.45 49.87
Rajan et al. rajan2022cross $\times$ A+V+T - - 56.10 9.96 41.58 84.13 60.39 63.95 44.59 0.00 24.26 2.69 1.76 35.40 48.78
T-ESFL liu2022mafw $\times$ A+V+T - - 61.89 1.10 7.69 85.90 - 71.87 62.17 0.00 36.00 0.00 0.00 31.00 50.29
T-MEP* zhang2023transformer $\times$ A+V+T 111 25 53.03 19.32 40.65 79.94 55.89 74.17 53.48 2.15 26.61 1.15 5.10 37.41 50.96
T-MEP zhang2023transformer $\times$ A+V+T 111 25 56.95 18.19 42.89 81.62 60.14 71.60 58.22 3.21 30.53 2.27 7.51 39.37 52.85
We first present the results of 11 single-labeled emotions on MAFW liu2022mafw in Table 1. Compared with state-of-the-art audio-visual methods, we observe that the proposed method outperforms them by large margins across three different model scales. For example, HiCMAE-B surpasses the best-performing supervised method T-MEP zhang2023transformer by +5.48% UAR and +5.02% WAR, setting a new record of 42.65% UAR and 56.17% WAR on this dataset. Besides, even the smallest model HiCMAE-T also achieves higher performance (+3.00% UAR and +2.26% WAR) than T-MEP, while having 3$\times$ fewer parameters and reducing more than 36% FLOPs. More surprisingly, although incorporating the textual information (i.e., descriptive sentences of affective behaviors) which is beneficial to emotion recognition on this dataset zhao2023dferclip ; foteinopoulou2023emoclip , T-MEP still lags behind the proposed method by a large margin. It also should be noted that, T-MEP employs strong pre-trained models (e.g., DeiT touvron2021training and RoBERTa liu2019roberta ) as unimodal encoders and shows degraded performance without using them. Therefore, these comparison results greatly demonstrate the powerful learning capacity of our method and the superiority of large-scale self-supervised pre-training over traditional supervised learning.

### v2

@@ 3.1.2 Encoder
Modality-Specific Encoder. As a general self-supervised pre-training framework, there can be many choices for the architecture of modality-specific encoders. For simplicity, we follow MAE [17] and VideoMAE [61] to employ standard Transformer [64] as audio and video encoders. Other efficient architectures such as LGI-Former [50] can be explored in future work. The audio and video encoders consist of $N_{s}$ Transformer layers. Each Transformer layer is mainly composed of Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN):
@@
where $\mathbf{E}^{0}_{m}=\mathbf{X}_{m}^{{}^{\prime\prime}}$, $m\in\{a,v\}$ denotes the audio or video modality, $j\in\{1,...,N_{s}\}$ is the layer index, and LN stands for layer normalization [64]. For MHSA in Eq. (1), it calculates dot-product attention to learn intrinsic dependency relationships in input audio/video tokens:
@@
where $\mathbf{W}^{Q}_{h}\in\mathbb{R}^{C\times d_{h}}$, $\mathbf{W}^{K}_{h}\in\mathbb{R}^{C\times d_{h}}$, $\mathbf{W}^{V}_{h}\in\mathbb{R}^{C\times d_{h}}$, $\mathbf{W}^{O}\in\mathbb{R}^{C\times C}$, $H$ is the number of attention heads, and $d_{h}=C/H$ is the feature dimension in each attention head. For FFN in Eq. (1), it comprises two linear projection layers with a GELU [66] activation function in between:
@@
Cross-Modal Fusion Encoder. After unimodal feature encoding, HiCMAE employs a cross-modal fusion encoder to capture meaningful interactions in audio-visual modalities and enable cross-modal reinforcement [51, 57]. We mainly utilize multi-head cross-attention (MHCA) to implement this module. MHCA shares the same spirit with MHSA. The main difference is that MHCA accepts two modalities as input, with one as the target modality and the other as the source modality. The formulation of MHCA is given as follows:
@@ 3.1.3 Decoder
After feature encoding, HiCMAE utilizes two modality-specific decoders for final masked audio-visual reconstruction. Previous studies have demonstrated that a small-capacity model is enough to reconstruct masked information [17, 61, 65]. Therefore, we follow them to employ a lightweight Transformer-based model as the decoder, whose number of layers is much less than that of the encoder, to largely reduce the computation burden during self-supervised pre-training.
@@
Hierarchical Skip Connections. As shown in Fig. 4, unlike previous audio-visual masked autoencoders which solely operate on encoder representation from the last layer and neglect explicit guidance on other layers [18, 63, 19, 20], HiCMAE incorporates hierarchical skip connections between intermediate encoder and decoder layers to explicitly steer encoder feature learning of different levels and assist the decoder to complete the task of masked audio-visual reconstruction. Specifically, HiCMAE adds an MHCA layer before each (except for the first) Transformer layer in the decoder. Given the tokens $\mathbf{D}^{k}_{m}$ ($m\in\{a,v\}$) output by the $k$th Transformer layer, the skip connection constructed by MHCA enables it to directly attend to the intermediate encoder representations $\mathbf{E}^{j}_{m}$ with different levels of information (from high-level coarse-grained to low-level fine-grained) and extract useful information to reinforce itself. Then, as usual, a standard Transformer layer is followed to exploit global self-attention to infer masked information. The whole process can be formulated as follows:
@@
where $\mathbf{X}_{m}$ ($m\in\{a,v\}$) denotes the original input, $\hat{\mathbf{X}}_{m}$ is the reconstructed data, $\mathcal{M}_{m}$ denotes the masked positions. Note that, for the visual modality, we follow [50] to reconstruct both spatial appearance and temporal motion information (i.e., frame difference signals).
@@ 3.2 Hierarchical Cross-Modal Contrastive Learning
To align with hierarchical skip connections in Section 3.1.3, we use the same $N_{c}$ selected encoder layers ($i.e.,\{q_{1},...,q_{N_{c}}\}$, $q_{i}\in\{1,...,N_{s}\}$) for HCMCL. Since HCMCL is conducted within a batch of samples, for convenience, we add an additional subscript $i$ to the notation of encoder tokens (i.e., $\mathbf{E}^{j}_{i,m}$) to indicate the sample index in this batch. Given a batch of $N$ audio tokens $\{\mathbf{E}^{j}_{1,a},...,\mathbf{E}^{j}_{N,a}\}$ and $N$ video tokens $\{\mathbf{E}^{j}_{1,v},...,\mathbf{E}^{j}_{N,v}\}$ from $j$th ($j\in\{q_{1},...,q_{N_{c}}\}$) encoder layer, we first perform global average pooling to obtain the sample-level features, i.e., $\mathbf{e}^{j}_{i,a}=\textrm{AvgPool}(\mathbf{E}^{j}_{i,a})$ and $\mathbf{e}^{j}_{i,v}=\textrm{AvgPool}(\mathbf{E}^{j}_{i,v})$. After that, we utilize symmetric InfoNCE loss [67] for HCMCL. For a batch of audio features $\mathbf{e}^{j}_{a}=\{\mathbf{e}^{j}_{1,a},...,\mathbf{e}^{j}_{N,a}\}$ and video features $\mathbf{e}^{j}_{v}=\{\mathbf{e}^{j}_{1,v},...,\mathbf{e}^{j}_{N,v}\}$, the symmetric InfoNCE loss maximizes the cosine similarity of $N$ paired audio-visual features (i.e., from the same sample) in the batch while minimizing the cosine similarity of features of $N(N-1)$ incorrect pairings (i.e., from different samples):
@@ 3.3 Overall Pre-training Loss
During self-supervised pre-training on massive unlabeled AVER data, the audio-visual emotion semantics are implicitly modeled by $\mathcal{L}$ in Eq. (11). Despite implicit modeling, the visualization analysis of masked audio-visual reconstruction in Section 4.6.1 shows that the emotion-related information (e.g., smiles in video frames and harmonics in the audio spectrogram) can be well restored by reasoning in limited visible contexts, indicating that HiCMAE can capture audio-visual emotion semantics via large-scale self-supervised pre-training.
@@ 3.4 Hierarchical Feature Fusion for Downstream Fine-tuning
After self-supervised pre-training, we discard the lightweight decoders and only use the encoders in HiCMAE for downstream fine-tuning on limited labeled AVER data. To benefit downstream emotion recognition tasks, we utilize both cross-modal and unimodal features from the encoders. The cross-modal features come from the cross-modal fusion encoder. We simply perform global average pooling to output tokens from the last fusion layer. The unimodal features come from modality-specific encoders. To fully exploit features of different levels, we use learnable weights to combine features from different audio/video encoder layers. Thus, the overall feature can be obtained as follows:
@@
After obtaining the overall feature $\mathbf{e}$, we simply use a linear layer to project it to get the final emotion prediction $\hat{\mathbf{y}}$. For the classification task, we utilize the classic cross-entropy loss for model fine-tuning:
$\mathcal{L}_{\textrm{CLS}}=-\sum_{k=1}^{K}{y_{k}\log{\hat{y}_{K}}},$ (13)
where $\hat{\mathbf{y}}=[\hat{y}_{1},...,\hat{y}_{K}]\in\mathbb{R}^{K}$ denotes the prediction, $\mathbf{y}=[y_{1},...,y_{K}]\in\mathbb{R}^{K}$ is the target, and $K$ is the number of emotion categories. For the regression task, we compute the mean square error between the target and the prediction:
@@
where $\hat{\mathbf{y}}\in\mathbb{R}^{D}$, $\hat{\mathbf{y}}\in\mathbb{R}^{D}$, and $D$ is the number of emotion dimensions. During downstream fine-tuning, the audio-visual emotion semantics are explicitly modeled by $\mathcal{L}_{\textrm{CLS}}$ in Eq. (13) or $\mathcal{L}_{\textrm{REG}}$ in Eq. (14).
@@ 4.1 Implementation Details
We pre-train HiCMAE on a very large audio-visual dataset VoxCeleb2 [27]. It contains more than one million video clips from over six thousand celebrities. VoxCeleb2 is split into a development set and a test set. In this paper, we only use its development set which has 1,092,009 video clips for self-supervised pre-training. For each video clip, we sample 16 consecutive frames with a temporal stride of 4 frames and follow [50] to extract $160\times 160$ patch in each frame to obtain the video input with a size of $16\times 160\times 160\times 3$. The corresponding audio waveform (2.56s) is also extracted and converted into a 128-dimensional log Mel filterbank feature sequence using a 25ms Hanning window with a hop length of 10ms [18], resulting in the audio spectrogram with a size of $256\times 128$. The loss weight in Eq. (11) is set to $\lambda=0.0025$. The temperature factor for contrastive learning is fixed to $\tau=0.07$. We pre-train HiCMAE for 100 epochs using four Nvidia Tesla V100 GPUs with a total batch size of 160 and a base learning rate of $43e-$ 4. It takes about five days to complete the pre-training. For downstream tasks, we fine-tune the pre-trained model for 50 or 100 epochs with a total batch size of 56 and a base learning rate of $31e-$ 3. During inference, we follow [50] to uniformly sample two clips from each video and calculate their mean score as the final prediction. Other hyper-parameters for pre-training and fine-tuning can refer to [61, 65, 18] for details.
@@ 4.2.1 Datasets
The basic dataset information is summarized in Table 1.
Table 1: Basic information of nine AVER datasets used in this paper.
Dataset Emotion Type Acquisition Condition #Subjects #Samples #Emotions Evaluation
MAFW[5] Categorical In-the-wild N/A 9,172 11 5-fold cross-validation (official)
Categorical In-the-wild N/A 8,996 43 5-fold cross-validation (official)
DFEW [29] Categorical In-the-wild N/A 11,697 7 5-fold cross-validation (official)
MER-MULTI [25] Categorical In-the-wild N/A 3,784 6 Official split
CREMA-D [28] Categorical Lab-controlled 91 7,442 6 5-fold cross-validation (subject-independent)
Categorical Lab-controlled 91 4,896 4 The last fold (subject-independent)
MSP-IMPROV [68] Categorical Lab-controlled 12 7,798 4 6-fold cross-validation (session-independent)
RAVDESS [69] Categorical Lab-controlled 24 1,440 8 6-fold cross-validation (subject-independent)
IEMOCAP [70] Categorical Lab-controlled 10 5,531 4 5-fold cross-validation (session-independent)
Werewolf-XL [23] Dimensional Lab-controlled 129 14,632 3 5-fold cross-validation (subject-independent)
AVCAffe [24] Dimensional In-the-wild 106 58,112 2 Official split
MAFW [5] is a large-scale multimodal compound in-the-wild affective dataset. It consists of 10,045 video clips annotated with 11 common emotions (including seven basic emotions, contempt, anxiety, helplessness, and disappointment). Each video clip is also accompanied by several textual sentences to describe the subject’s affective behaviors. The dataset provides an 11-class single-labeled set (9,172 video clips) and a 43-class compound set (8,996 video clips). For model evaluation, we follow the original paper to adopt a 5-fold cross-validation protocol.
DFEW [29] comprises 16,372 video clips which are extracted from over 1,500 high-definition movies. This dataset presents several challenging characteristics, such as extreme illumination and occlusion. The video clips are annotated with seven basic emotions (i.e., happy, sad, neutral, anger, surprise, disgust, and fear). To align with previous work [29, 50, 9], we perform 5-fold cross-validation on 11,697 single-labeled clips for evaluation.
MER-MULTI [25] provides 3,373 training video clips originating from Chinese TV series and movies. The dataset is annotated with six emotions, including neutral, anger, happiness, sadness, worry, and surprise. We follow the original paper to conduct 5-fold cross-validation on 3,373 video clips for hyperparameter tuning and evaluate the model on a held-out test set with 411 video clips.
@@ 4.2.2 Results on MAFW
Wav2Vec2.0 [47] A 95 18 59.01 9.39 26.08 31.47 32.04 46.52 9.91 1.69 12.23 3.05 6.04 21.59 29.69
HuBERT [48] A 95 18 54.97 15.49 31.20 28.64 36.88 58.39 12.52 2.54 12.55 5.34 16.48 25.00 32.60
WavLM-Plus [71] A 95 18 55.62 17.21 40.48 36.65 36.53 57.44 11.12 2.12 11.35 9.54 11.54 26.33 34.07
@@
ResNet-18 [72] $\times$ V 11 - 45.02 9.25 22.51 70.69 35.94 52.25 39.04 0.00 6.67 0.00 0.00 25.58 36.65
ViT [73] $\times$ V - - 46.03 18.18 27.49 76.89 50.70 68.19 45.13 1.27 18.93 1.53 1.65 32.36 45.04
C3D [43] $\times$ V 78 39 51.47 10.66 24.66 70.64 43.81 55.04 46.61 1.68 24.34 5.73 4.93 31.17 42.25
ResNet-18+LSTM [5] $\times$ V - - 46.25 4.70 25.56 68.92 44.99 51.91 45.88 1.69 15.75 1.53 1.65 28.08 39.38
ViT+LSTM [5] $\times$ V - - 42.42 14.58 35.69 76.25 54.48 68.87 41.01 0.00 24.40 0.00 1.65 32.67 45.56
C3D+LSTM [5] $\times$ V - - 54.91 0.47 9.00 73.43 41.39 64.92 58.43 0.00 24.62 0.00 0.00 29.75 43.76
Former-DFER [74] $\times$ V 18 9 58.23 11.45 31.29 75.06 43.07 63.81 46.02 0.42 26.22 2.88 2.25 32.79 45.31
T-ESFL [5] $\times$ V - - 62.70 2.51 29.90 83.82 61.16 67.98 48.50 0.00 9.52 0.00 0.00 33.28 48.18
T-MEP [9] $\times$ V 5 6 52.91 17.41 28.01 80.79 49.42 58.73 49.54 0.00 26.18 2.25 3.56 33.53 47.53
DFER-CLIP [75] V 153 92 - - - - - - - - - - - 39.89 52.55
SVFAP [49] V 78 44 64.60 25.20 35.68 82.77 57.12 70.41 58.58 8.05 32.42 8.40 9.89 41.19 54.28
MAE-DFER [50] V 85 50 67.77 25.35 34.88 77.13 58.26 71.09 57.46 8.90 33.08 11.83 12.09 41.62 54.31
@@
ResNet-18+LSTM [5] $\times$ A+V - - 54.47 11.89 7.07 82.73 54.85 55.06 39.35 0.00 15.99 0.39 0.00 29.26 42.69
C3D+LSTM [5] $\times$ A+V - - 62.47 3.17 15.74 77.30 42.20 65.30 42.67 0.00 19.14 0.00 0.00 30.47 44.15
AMH [76] $\times$ A+V - - 51.73 18.68 28.13 79.14 52.55 52.26 46.29 0.26 29.62 1.74 2.39 32.98 48.83
T-ESFL [5] $\times$ A+V - - 60.73 1.26 21.40 80.31 58.24 75.31 53.23 0.00 14.93 0.00 0.00 33.35 48.70
T-MEP* [9] $\times$ A+V 61 22 54.98 22.11 32.23 82.79 50.90 62.50 49.93 0.87 29.27 8.09 6.70 36.40 48.17
T-MEP [9] $\times$ A+V 61 22 57.04 24.85 36.09 78.96 50.83 61.85 51.28 1.29 38.47 6.46 1.70 37.17 51.15
@@
ResNet18+MDRE [77] $\times$ A+V+T - - 45.59 9.35 24.30 76.31 51.10 74.87 28.82 2.08 30.99 0.00 0.00 31.22 48.33
AMH [76] $\times$ A+V+T - - 54.91 19.41 30.01 82.79 51.42 60.73 51.54 0.00 28.18 0.00 0.00 34.45 49.87
Rajan et al. [78] $\times$ A+V+T - - 56.10 9.96 41.58 84.13 60.39 63.95 44.59 0.00 24.26 2.69 1.76 35.40 48.78
T-ESFL [5] $\times$ A+V+T - - 61.89 1.10 7.69 85.90 - 71.87 62.17 0.00 36.00 0.00 0.00 31.00 50.29
T-MEP* [9] $\times$ A+V+T 111 25 53.03 19.32 40.65 79.94 55.89 74.17 53.48 2.15 26.61 1.15 5.10 37.41 50.96
T-MEP [9] $\times$ A+V+T 111 25 56.95 18.19 42.89 81.62 60.14 71.60 58.22 3.21 30.53 2.27 7.51 39.37 52.85
We first present the results of 11 single-labeled emotions on MAFW [5] in Table 2. Compared with state-of-the-art audio-visual methods, we observe that the proposed method outperforms them by large margins across three different model scales. For example, HiCMAE-B surpasses the best-performing supervised method T-MEP [9] by +5.48% UAR and +5.02% WAR, setting a new record of 42.65% UAR and 56.17% WAR on this dataset. Besides, even the smallest model HiCMAE-T also achieves higher performance (+3.00% UAR and +2.26% WAR) than T-MEP, while having 3$\times$ fewer parameters and reducing more than 36% FLOPs. More surprisingly, although incorporating the textual information (i.e., descriptive sentences of affective behaviors) which is beneficial to emotion recognition on this dataset [75, 79], T-MEP still lags behind the proposed method by a large margin. It also should be noted that T-MEP employs strong pre-trained models (e.g., DeiT [80] and RoBERTa [81]) as unimodal encoders and shows degraded performance without using them. Therefore, these comparison results greatly demonstrate the powerful learning capacity of our method and the superiority of large-scale self-supervised pre-training over traditional supervised learning.